### v1

Abstract
We present a framework, DISORF, to enable online 3D reconstruction and visualization of scenes captured by resource-constrained mobile robots and edge devices. To address the limited compute capabilities of edge devices and potentially limited network availability, we design a framework that efficiently distributes computation between the edge device and remote server. We leverage on-device SLAM systems to generate posed keyframes and transmit them to remote servers that can perform high quality 3D reconstruction and visualization at runtime by leveraging NeRF models. We identify a key challenge with online NeRF training where naive image sampling strategies can lead to significant degradation in rendering quality. We propose a novel shifted exponential frame sampling method that addresses this challenge for online NeRF training. We demonstrate the effectiveness of our framework in enabling high-quality real-time reconstruction and visualization of unknown scenes as they are captured and streamed from cameras in mobile robots and edge devices.
I Introduction
Online 3D reconstruction to learn a representation of a scene at real-time—where RGB images are continuously captured and used to optimize a 3D model that can be rendered and visualized at real-time—holds immense potential in various domains. For example, mobile robots and embodied devices (e.g., drones) navigating a previously unseen environment with the ability to construct and visualize a 3D model of the environment on the fly, offer opportunities for enhanced navigation, scene understanding, and interactive exploration of the environment. Online 3D scene reconstruction has been extensively studied with various methods to represent geometry and appearance based on voxel or point representations [1, 2, 3]. Recently, implicit neural representation methods such as neural radiance fields (NeRFs) [4] have emerged as a promising approach to representing complex 3D scenes with the capability of photorealistic 3D scene rendering and visualization.
NeRF methods use multi-view posed images to learn a neural implicit function (usually an MLP) optimized through differentiable volumetric rendering. NeRF is commonly used as an offline 3D scene representation method as its training process requires sampling over multi-view images with well-calibrated camera parameters. However, to make NeRF compatible with an incremental online learning process, the camera poses of captured frames need to be estimated on the fly. This can be done by leveraging the tracking module in real-time simultaneous localization and mapping (SLAM) systems typically used in mobile robots. The integration of NeRF and SLAM is a great opportunity to unlock NeRF’s capability for learning and visualizing 3D scene representations online. Recent research has also investigated the potential of leveraging SLAM for online 3D reconstruction. However, there are several challenges that need to be addressed to make NeRF-based online 3D scene representation feasible in mobile robots.
Figure 1: The setup and rendering results for online and offline NeRF training: For offline NeRF training (top), images from all viewpoints are available ahead of time. For online NeRF training (bottom), however, the model is continuously trained as new images are streamed.
First, mobile and embodied robots or edge devices are typically highly resource-constrained, being unable to support powerful GPUs and large memory capacities. Thus, performing the computationally expensive NeRF training on resource-constrained edge devices is impractical. Although recent advances have seen great improvements in accelerating NeRF training and rendering speed [5, 6, 7], the substantial computing requirement for efficient NeRF training cannot be satisfied on resource-constrained edge computing platforms. For example, a powerful edge GPU device like Jetson Xavier NX [8] still takes over 14 times longer [9] than an RTX3090 GPU to train Instant-NGP, currently one of the most efficient NeRF models. Distribution of compute to a remote server that is provisioned with more powerful compute resources is a promising approach to enable expensive online NeRF training and rendering. Recent work [10] has developed a framework for transmitting image streams from edge cameras and optimizing them offline on a remote server. However, this approach relies solely on remote computation, neglecting the potential benefits of using mobile robots’ visual odometry and localization capabilities for pose estimation. Additionally, it assumes a constant and sufficient network bandwidth, limiting its applicability to specific domains and use cases.
Second, a critical challenge in enabling high-quality online 3D reconstruction is the sampling strategy for online NeRF training. NeRF training requires sampling a batch of rays/pixels from captured images of the scene being represented. The commonly used sampling method is random uniform sampling, which uniformly samples $N$ rays from all existing images for each training iteration. This approach is effective for offline training, however, leads to sub-optimal rendering with online training as shwon in Fig. 1. To illustrate this challenge, let’s consider an example comparing online and offline NeRF training scenarios. In both cases, we train the NeRFacto [11] model using keyframes generated by on-device ORB-SLAM2 [12] from a 1-minute Replica RGB video stream. As depicted in Fig. 2, online NeRF training with uniform ray sampling results in significantly worse rendering quality, with scene objects appearing blurred. This quality drop can be partly explained by the unbalanced frame sampling distribution. As shown in Fig. 3, offline training samples a roughly equal number of rays/pixels from each training frame, whereas online training ends up sampling more from the earlier frames and less from the recent frames overall. This is because, as the training process continues while the images are being continuously streamed, at any given time, the earlier images can be sampled in more iterations. Since the more recent frames are less sampled, objects in these frames may not be sufficiently sampled in the online training process to optimize their shape and appearance, thus causing the lower rendering quality.
Our goal is to enable online 3D reconstruction and visualization of environments/scenes from mobile robots and edge devices by addressing the challenges mentioned above. To achieve this, we introduce DISORF, a novel framework that enables online 3D reconstruction with NeRF by distributing the computational tasks between the edge devices and a remote server. With DISORF, we leverage on-device SLAM for pose estimation and only transmit keyframes to the remote server for processing. This approach effectively reduces the reliance on network bandwidth for high-quality reconstruction. The resource-intensive NeRF computations are performed on powerful servers, essential for online NeRF training and rendering.
We explore input sampling strategies for online 3D reconstruction and find that giving greater weight to recent frames during each iteration visibly enhances reconstruction quality. Building on this insight, we introduce a shifted exponential sampling weight function. This function dynamically focuses on recently received frames during training, mitigating the issue of inadequate samples for recent frames.
Figure 2: The rendering results from different NeRF training strategies. The synthesized view is from a camera trajectory on Replica office4. All methods are trained for the same number of training iterations.
Figure 3: The total number of pixels from each frame being sampled after training on Replica office4’s one-minute keyframe stream.
II Related Work
II-A Online 3D Representation
Unlike offline reconstruction methods that are able to access all the frames for iterative global optimizations [13, 14], the online 3D reconstruction task requires on-the-fly camera pose estimation and reconstruction. Therefore, most online 3D reconstruction is closely integrated with SLAM systems. These dense 3D reconstruction methods [15, 16, 17, 18, 19] typically use RGB-Depth sensors as input to incrementally reconstruct scene geometry using either the signed distance field (SDF) or occupancy estimated with voxels as 3D scene representations and some approaches use explicit representations such as surfels [20, 3, 21]. With the emergence of neural representations [22, 23], recent works have leveraged neural networks for online 3D reconstruction [24, 25, 26]. Recent neural reconstruction methods also enable promising online 3D reconstruction results from monocular video streams [27, 28, 29]. These prior works aim to enable high-quality 3D surface or geometry reconstruction and do not address the challenges of enabling high-quality online 3D reconstruction and visualization on resource-constrained mobile robots and edge devices.
II-B Neural Radiance Fields and Robotics Applications
NeRF [4] is an implicit neural scene representation that enables photorealistic view synthesis given a set of multiview posed images. The key idea is to use a neural network (e.g., an MLP) to represent implicit fields (e.g., volume density and color) that are used for traditional volumetric rendering and is optimized using a photometric loss. The training process of the initial NeRF architecture [4] is compute-intensive, requiring hours or even days, and has high rendering latencies. Recent advances [5, 6, 7] have significantly accelerated NeRF’s training and rendering speed, requiring only minutes or even seconds of training time. Our online NeRF training platform is built on Nerfstudio [11] that enable high-speed offline NeRF training and rendering.
NeRF’s promising capabilities have also been leveraged for various robotics tasks. Evo-NeRF [30] trains NeRFs online while grasping using a robotic arm. This is primarily focused on the high-speed training of NeRFs given an evolving scene as opposed to achieving high-quality online 3D reconstruction. Existing works [31, 32, 33, 34] leverage NeRF for scene representation to facilitate downstream tasks such as trajectory planning, training a control policy for robot motion, etc. However, these works use a pre-trained implicit model for various downstream tasks, while our work focuses on online training. Our framework could be potentially used in these scenarios to enable online NeRF training.
II-C Incremental Learning for NeRF
Incremental or continual learning is a task in which a machine learning model has to learn new knowledge from a training dataset which is continually expanded as new data is gathered over time. Incremental learning often encounters two significant challenges: forgetting, where the model may lose previously learned knowledge when exposed to new data [35], and adaptation, which involves efficiently incorporating new information without compromising the existing knowledge [36]. Various approaches have been proposed for incremental learning [37] to address these challenges. The online NeRF training can be treated as a replay-based incremental learning approach [38, 39]. There are some existing works that perform incremental training of implicit neural representations in SLAM systems for dense surface reconstruction [40, 41, 42]. For incremental training of NeRFs, these neural SLAM methods often manually pre-define a fixed portion of pixels to sample from recent frames and the remainder from earlier frames in each training iteration. In contrast, our sampling method offers a smoother transition from concentrated to uniform sampling as the training iteration progresses and the number of training frames increases, which enhances its adaptability across diverse scenes and scenarios. We compares against these approach in Sec. IV and demonstrate visibly better rendering quality.
III Method
$\alpha$=1,$\beta$=0
[…]
Iteration $S=124$
Iteration $S=128$
Iteration $S=132$
Iteration $S=141$
Figure 4: Sampling count distribution, illustrating the effect of our shifted exponential sampling weight functions with varying $\alpha$ and $\beta$ values at different training iterations. The horizontal axis represents the frame index, while the vertical axis indicates the number of sampled rays. Each iteration contains a total of 8192 samples. Notably, we emphasize the fluctuations in sampling counts for frame #30, received at iteration 124.
We now describe our proposed framework, DISORF, to enable online 3D scene reconstruction and visualization from images streamed from a mobile robot or edge device. We first describe our distributed framework that enables offloading some computing to a remote server to support the expensive computations required for online 3D reconstruction. We then describe our improved sampling strategy suited for online NeRF training. Since the pixel/ray sampling method is independent of the NeRF model itself, the proposed method is also applicable to various radiance field models.
III-A Online NeRF Training Framework for Robotics
Figure 5: Pipeline of DISORF that enables the distribution of computation between the mobile robot/edge device and a remote server. The camera on the mobile robot captures RGB images that are processed by the SLAM module on the edge processor. The generated keyframes and pose information are continually streamed over the network to a central server. The more compute-intensive NeRF training is performed on remote servers with powerful GPUs.
As illustrated in Fig. 5, our framework enables distributing computation between a local end (mobile robot) and a remote server connected across the network. Our framework comprises an interface at the local end that takes the camera image as input and produces keyframes and estimated poses as output. The SLAM module used to generate the keyframes and estimated poses is configurable in the framework. DISORF then transmits the keyframes and estimated poses using a network connection through ROS [43] interface to a remote process running on a server with sufficient compute resources for NeRF training. The local end of our framework encapsulates both the image and pose into a network packet. This packet is then transmitted via a publisher through the network and forwarded to the remote end. Given the limited amount of data that is transmitted with our framework (i.e., posed keyframes), our framework is highly resilient to the limited availability of network bandwidth. Compared to the solution that sends all the video frames to the remote server, our keyframe-based transmission has over 10x reduction in network bandwidth usage (e.g., 5.2MBps vs 358.4KBps, tested on Replica scan). Our evaluation framework can be used on mobile robots with mobile GPUs like NVIDIA Jetson series edge processors. The remote end can be interfaced with any 3D reconstruction training module. In our experiments, we use Nerfstudio [11] to perform training and rendering tasks as the robot explores the scene. Its subscriber receives the packet from the local end and stores it in the database. The model being learned online can be visualized in real-time at the remote end.
III-B Modeling the Keyframe Stream
Under the online training setting, each new keyframe $I_{i}$ is sequentially sent to the NeRF training server at time $T_{i}$. The time interval between two adjacent keyframes $\Delta T_{i}=T_{i+1}-T_{i}$ varies due to factors like camera or picture motion, system computation, and network fluctuations. We can loosely model the arrival of new keyframes as a poisson point process (PPP). By definition of the Poisson point process, the time interval $\Delta T_{i}$ between successive keyframes follows an exponential distribution:
$\Delta T\sim Exp(\lambda),\quad f(x,\lambda)=\lambda\exp(-\lambda x)$ (1)
where $f(x,\lambda)$ is PDF function, $\lambda$ is the expected rate of arrival of a new keyframe. This modeling of the keyframe stream gives us insights for designing a new sampling method that emphasizes recent frames while considering the keyframe arrival rate.
III-C Sampling with Shifted Exponential Distribution
The sampling problem we deal with is determining which keyframes we should sample rays/pixels from for each training iteration, given the limited number of total iterations for online training. As discussed in Section I, the naive uniform sampling approach does not sufficiently sample the later frames resulting in a rendering quality drop. Thus, we need a sampling method that generates more samples from the more recently received frames during training.
Before presenting our method, we first define how we represent time for NeRF training. Since the sampling is performed once per training iteration, and these iterations roughly take the same amount of time, we use the number of iterations to denote elapsed time. For instance, the timestamp $T_{i}$ of keyframe $I_{i}$ corresponds to the $T_{i}$-th iteration when the NeRF training routine first encounters $I_{i}$.
We seek to define a function $f$ that maps each keyframe to a sampling weight. The earlier frames that are far behind the current iteration should have lower weights and perform a more uniform sampling, while the newer frames should have higher weights. To ensure this function remains unaffected by the continuous increase in training iterations, we use relative time intervals as input for $f$, instead of absolute timestamps. Specifically, the time interval $D_{i}$ for a keyframe with timestamp $T_{i}$ at iteration $S$ is calculated as $D_{i}=S-T_{i}$, ensuring that newer frames consistently have lower $D_{i}$. Thus, the function $f(D)$ should be decreasing over time.
Inspired by the exponential distribution of time intervals between two adjacent keyframes (Eq. 1), we utilize the properties of an exponential distribution to define a sampling weight function [^1] The sampling weights $W$ will be normalized for the PDF sampling.:
$W^{*}_{i}=f^{*}(D_{i})=\exp(-\lambda D_{i})$ (2)
However, simply using this function would severely downweight the earlier frames. Because of the exponential decrease, the earlier frames will have sampling weights very close to zero and almost all the samples will come from the recent frames. The lack of access to the earlier frames could cause the forgetting issue [35] in NeRF training which could further cause the overall rendering quality drop [40, 44].
To address this problem, we propose to add an offset term $\beta/N_{S}$ to ensure the earlier frames still have certain sampling weights to be almost uniformly sampled:
$W_{i}=f(D_{i})=\exp(-\alpha\lambda D_{i})+\beta/N_{S}$ (3)
where $N_{S}$ is the number of available keyframes at iteration $S$; $\alpha$ and $\beta$ are two hyperparameters: $\alpha$ scales the average keyframe rate to further control the decrease rate of $f(D)$, $\beta$ can control the ratio of the rays to be sampled from the most recent frames (e.g., $T_{N_{S}}=S$), because the portion of rays being sampled from frame $I_{N_{S}}$, given a sufficiently large $N_{S}$, can be roughly approximated as
$p_{N_{S}}=\frac{W_{N_{S}}}{\sum_{i=1}^{N_{S}}W_{i}}=\frac{1+\beta/N_{S}}{\beta+\sum_{i=1}^{N_{S}}\exp(-\alpha\lambda D_{i})}\approx\frac{1}{\beta+2}$ (4)
Fig. 4 demonstrates the sampling results with varied $\alpha$ and $\beta$ values. Our proposed sampling function can gradually decrease the sampling weight of a newly received frame as training proceeds. With the help of the $\beta$ term, the sampling for earlier frames is similar to uniform sampling, providing sufficient training samples to avoid the forgetting issue.
TABLE I: Quantitative comparison on Replica scenes.
Method
off-0
off-1
off-2
off-3
off-4
rm-0
rm-1
rm-2
Avg.
PSNR normal-$\uparrow$
offline 34.31 33.52 27.78 27.61 29.44 26.80 28.16 29.42 29.63
uniform
33.23
32.42
26.88
27.09
28.06
25.96
27.47
29.07
28.77
imap frame
33.40
32.77
27.05
27.23
28.71
26.21
27.71
28.88
28.99
exp
33.51
32.77
27.12
27.24
28.51
26.27
27.67
29.42
29.06
loss only
33.75
33.51
27.32
27.02
28.27
26.30
27.81
29.09
29.13
imap loss 33.89 33.54 27.84 27.65 29.17 26.54 27.90 29.64 29.52
exp loss 34.43 33.71 27.82 27.88 29.23 26.63 28.09 29.79 29.70
SSIM normal-$\uparrow$
offline 0.921 0.901 0.882 0.873 0.901 0.776 0.819
0.879
0.869
uniform
0.905
0.885
0.876
0.862
0.890
0.752
0.815
0.880
0.858
imap frame
0.908
0.889
0.875
0.865
0.895
0.758
0.816
0.877
0.860
exp
0.909
0.884
0.879
0.867
0.894
0.760
0.816 0.883
0.861
loss only 0.913 0.899 0.880
0.863
0.891
0.760
0.819
0.879
0.863
imap loss 0.913
0.897
0.880 0.869 0.896 0.766 0.816 0.884
0.865
exp loss 0.919 0.901 0.883 0.872 0.898 0.768 0.821 0.887 0.869
LPIPS normal-$\downarrow$
offline 0.176 0.212
0.268
0.238 0.260 0.323 0.301 0.227 0.251
uniform
0.220
0.253
0.277
0.259
0.289
0.383
0.316
0.248
0.281
imap frame
0.223
0.232
0.278
0.259
0.275
0.362
0.313
0.247
0.274
exp
0.206
0.240
0.267 0.251
0.274
0.360
0.318
0.235
0.269
loss only
0.204
0.211 0.261
0.270
0.286
0.369
0.308
0.233
0.268
imap loss 0.202
0.213
0.262
0.253
0.273 0.356 0.305
0.238
0.263
exp loss 0.187 0.211
0.268
0.240 0.268 0.359 0.307 0.228 0.259
III-D Loss-Guided Active Ray Sampling
Similar to prior works that actively sample more rays/pixels on the regions with higher training loss [40, 5], our proposed frame sampling method in Sec. III-C can also be incorporated with a loss-guided active ray sampling method to improve the performance of the trained model further. The loss-guided active ray sampling method typically involves tracking the spatial loss distribution for each frame during training iterations. To achieve this, we divide each image into an $M\times M$ grid, where each grid patch maintains a running average L1 photometric loss of sampled training rays or pixels within that patch.
To incorporate loss active ray sampling with our shifted exponential frame sampling, we split the sampling into two parts. Suppose we need to sample $B$ rays for one iteration, we first use the sampling weight function (Eq. 4) to calculate the portion of frames $p_{r}$ that should be sampled from the frames that are received within the last $K$ training iterations. We then uniformly sample $B_{r}=\lfloor p_{r}B\rfloor$ rays from those recent frames according to their sampling weight. For the remaining $B-B_{r}$ rays, we sample them based on the patch loss distribution across all the rest unsampled frames.
Note that this loss active sampling is not well suited for real-captured outdoor scenes that contain significant background noise and many moving objects. In such cases, outliers often contribute to higher loss values, but optimizing these regions does not necessarily improve the overall model quality. In fact, it can degrade model performance by utilizing valuable sampling resources inefficiently. Hence, we simply use uniform ray sampling for noisy outdoor scenes.
IV Experiments
IV-A Implementation Details
To deploy our online NeRF framework, we use a desktop machine with an RTX4090 GPU as the remote server for NeRF training and a Jetson Xavier NX as the local edge processor, meeting SLAM computation needs efficiently with low energy consumption. The local end and remote server are connected through a WiFi network. We choose NeRFacto [11], a NeRF model that integrates many recent advancements and performance optimizations, as the default model in the experiments. We enable NeRFacto’s pose refinement feature [45] to make it more robust to less accurate camera poses from real-time tracking. We leverage a customized ORB-SLAM2 [12] as the on-device SLAM for the local end.
It is challenging to fairly compare different online approaches due to the impact of the randomness of the data generated by the local end and transmitted across the network. We therefore implement a simulation module on the remote end, which replays the keyframe stream logged by real-time SLAM systems. This approach guarantees that the keyframe stream exposed to the NeRF training process remains consistent across different training strategies for an accurate and fair comparison. The offline NeRF training in our experiments also uses the same set of keyframes but with full access to all the keyframes throughout the training. For all the evaluated online NeRF training strategies, we ensure that at any specific training iteration, the same number of keyframes is presented across different strategies. This is achieved by replaying a keyframe-iteration log, which is tracked by running the naive online NeRF training without any computation interference on the remote end. The online NeRF training stops upon receipt of all keyframes, with only a few additional training iterations (shorter than the average keyframe time interval). This tracked number of training iterations is then applied across all the evaluated training strategies. We evaluate scenes with PSNR, SSIM, and LPIPS metrics on the uniformly sampled frames along the camera trajectory to quantitatively evaluate the rendering quality.
TABLE II: Quantitative comparison on TnT scenes.
Scene Barn Train Truck
Method PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
offline 25.30 0.721 0.290 23.54 0.725 0.232 24.78 0.791 0.168
uniform
23.23
0.627
0.418
20.41
0.554
0.451
23.27
0.723
0.260
imap frame 23.91 0.660 0.368 21.40 0.603 0.378 23.59 0.737 0.232
exp 24.22 0.677 0.345 21.81 0.624 0.352 23.74 0.739 0.236
IV-B Datasets
We evaluate our method using the Replica [46] and Tanks and Temples [47] datasets. Replica contains various synthetic small-scale indoor scenes, we use the same camera trajectory of around 1-minute length from iMAP [40] for our online NeRF learning. The tanks and temples (TnT) dataset has a collection of high-resolution real-captured video recordings (3-7 minutes) of various indoor and outdoor scenes. We pick a small subset of outdoor scenes to showcase the capability of our method for challenging unbounded outdoor scenes. We utilize ORB-SLAM2 with downscaled frames (640x360) and a reduced frame rate (10 FPS) for more stable tracking performance on the TnT video input. We use RGB video streams of scenes for all evaluated methods.
office2
office3
room1
[…]
offline
uniform
imap frame
exp
imap loss
exp loss
Figure 6: Visual comparison of synthesized views from Replica scenes.
Barn
Truck
[…]
offline
uniform
imap frame
exp
Figure 7: Visual comparison of NeRF-synthesized TnT scenes.
IV-C Evaluation
We first include offline NeRF training (“offline”) and online NeRF training with naive uniform sampling (“uniform”) as two fundamental baselines for our comparison. We then compare the methods with the keyframe sampling strategies. We implement a keyframe sampling method similar to iMAP (“imap frame”) that always samples $1/5$ rays of training batch from the most recent keyframe and uniformly samples the rest of rays from remaining keyframes. For indoor scenes from Replica, we additionally compare methods that leverage the loss-guided active sampling method discussed in Sec. III-D. The first baseline is the pure loss active sampling (“loss only”) which utilizes patch-wise loss distribution to sample rays. We also combine “imap frame” with loss active sampling to create a stronger iMAP baseline (“imap loss”). We use “exp” to denote our shifted exponential sampling method, and “exp loss” to denote “exp” with loss active sampling. We use $\alpha=2$, $\beta=4$ for our sampling methods.
IV-D Result Comparison
Replica. In Table I and Fig. 6, we present quantitative and qualitative results for the Replica dataset, respectively. These results show that our shifted exponential sampling method improves the quality of renderings across various scenes. Our frame sampling method (“exp”) consistently does better than the naive uniform sampling and iMAP’s frame sampling. Furthermore, when combined with the loss-guided active sampling method, our method “exp loss” even outperforms the offline baseline on some replica scenes. The visual results in Fig. 6 demonstrate the noticeable quality disparities among different online training methods.
Tanks and Temples. We evaluate three complex outdoor scenes with results shown in Table II and Fig. 7. Our frame sampling method still consistently outperforms the naive uniform sampling method and iMAP frame sampling method on most metrics.
We conclude from these results that our shifted exponential sampling can effectively improve the rendering and visualization quality for online NeRF training from a streaming edge device.
IV-E Evaluation with Gaussian Splatting
To further demonstrate the applicability and effectiveness of our sampling method, we apply our sampling method to 3D Gaussian Splatting (3DGS) [48], a recently introduced point-based 3D representation method, for online 3D scene reconstruction. We evaluate the online 3DGS training on two indoor scenes from TnT dataset (with COLMAP [13] poses for this experiment). In our setting, 3DGS is initialized with 10K random points. The results in Table III indicate that our frame sampling method continues to enhance rendering quality, even when working with a substantially different 3D model. Conversely, iMAP’s frame sampling approach performs even worse than the basic uniform sampling method.
TABLE III: Metrics for online 3DGS training on TnT scenes.
Scene Church Meetingroom
Method PSNR SSIM LPIPS PSNR SSIM LPIPS
offline 23.33 0.812 0.251 26.88 0.902 0.162
uniform 22.49 0.790 0.276 26.05 0.884 0.187
imap frame
22.15
0.779
0.279
25.74
0.880
0.190
exp 22.55 0.788 0.273 26.24 0.888 0.183
V Conclusion
We introduce a distributed framework, DISORF, to enable online 3D reconstruction and visualization of scenes that are captured from cameras on resource-constrained mobile robots. DISORF leverages the posed keyframes computed by on-device SLAM and higher computational capabilities by leveraging a remote server over the network to enable online NeRF training. We observe and address a key challenge with naive online NeRF training compared to the offline training is the unbalanced frame sampling during training that leads to significant loss in quality. The experimental results on various datasets and models demonstrate the effectiveness and applicability of our proposed sampling method. We believe our framework has the potential to enable more use cases and downstream real-time robotics tasks by leveraging the high-quality online visualization and 3D representation.